# LieRE: Lie Rotational Positional Encodings

### v4

@@ B.6 Initialization Sensitivity
Table 7: Paired Z-test between $2\pi$ and $1$ initialization for LieRE8 and RoPE-Mixed.
@@ Appendix C Python implementation of LieRE rotation matrix computation
basis_raw_params = nn.Parameter( torch.rand( input_dimensionality, head_dim, head_dim, ) * 2 * math.pi # optional, inspired from RoPE-Mixed paper ) upper_triangle = ( torch.triu(basis_raw_params, diagonal=1) ) skew_bases = upper_triangle - torch.transpose(upper_triangle, -1, -2) in_basis_positions = ( positions.reshape(list(positions.shape) + [1] * 2) * skew_bases ) rotation_log = torch.sum(in_basis_positions, dim=-3) rotation = torch.matrix_exp(rotation_log.to(dtype=torch.float32)) rotation = rotation.to(dtype=positions.dtype)
@@ Appendix D Compute efficiency
We demonstrate that LieRE-based transformer achieves comparable performance to the Absolute Position Embedding baseline (DeiT III (Touvron et al., 2022)) on CIFAR-100 with fewer training epochs. This represents a notable advancement over recent methods such as VisionLlama and RoPE-Mixed (Chu et al., 2024; Heo et al., 2024). Figure 9 illustrates that LieRE enables a 3.9$\times$ reduction in training compute while maintaining the accuracy achieved by absolute position encodings after 200 epochs.

### v5

@@ B.6 Initialization Sensitivity
Table 7: Paired Z-test between $2\pi$ and 11 initialization for LieRE8 and RoPE-Mixed.
@@ Appendix C Python implementation of LieRE rotation matrix computation
basis_raw_params = $n$.Parameter( torch.rand( input_dimensionality, head_dim, head_dim, ) * 2 * math.pi # optional, inspired from RoPE-Mixed paper ) upper_triangle = ( torch.triu(basis_raw_params, diagonal=1) ) skew_bases = upper_triangle - torch.transpose(upper_triangle, -1, -2) in_basis_positions = ( positions.reshape(list(positions.shape) + [1] * 2) * skew_bases ) rotation_log = torch.sum(in_basis_positions, dim=-3) rotation = torch.matrix_exp(rotation_log.to(dtype=torch.float32)) rotation = rotation.to(dtype=positions.dtype)
@@ Appendix D Compute efficiency
[…]